OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = x^2 + 8x + 12. (-4, 12) (-4, -4) (0, -6)

OpenStudy (jadeishere):

-4 and -4

OpenStudy (anonymous):

Find the distance between the points (-4, -5) and (3, -1).

OpenStudy (anonymous):

\[\sqrt{65} or \sqrt{37} or \sqrt{17}\] choices for previous question